ABC Company used linear regression theory to develop the slope and intercept values for a forecast equation in Y (sales volume) and X (number of customers) using the data below (sales in thousands of dollars; customers in hundreds). Forecast the sales volume when the when the number of customers is 22. a = –0.158; b = 0.131

USTOMERS(X)

10

12

16

15

14

17

20

SALE VOLUME(Y)

1.0

1.4

1.9

2.0

1.8

2.1

2.3
Answer #1

Given, Regression equation Y = a+bX

Y = -0.158 + 0.131 X

When the customers are 22, X=22

Sales Y = -0.158+ 0.131*22 = 2.724

Sales at x=22 is Y = 2.724